OpenStudy (anonymous):

cos(2x) less than or equal to 0 solve the inequality

myininaya (myininaya):

\[\cos(2x) \le 0\] first lets think about \[\cos(x) \le 0\] but even then lets first think about \[\cos(x)=0\] \[\cos(x)=0=> x=n \pi, n \in \mathbb{Z}\] so this means \[\cos(2x)=0 when 2x=n \pi => x=\frac{n \pi}{2}\] now let's think about \[\cos(x) <0\] on the interval \[(0,2\pi)\] cos(x) is negative on the interval \[(\frac{\pi}{2}, \frac{3\pi}{2})\] so cos(2x) is negative on the interval \[(\frac{1}{2} \cdot \frac{\pi}{2}, \frac{1}{2} \cdot \frac{3\pi}{2})=(\frac{\pi}{4}, \frac{3\pi}{4})\] but now we got to think for all real numbers that was just on the interval 0 to 2pi so if we go around circle again and we look at the interval 2pi to 4pi we have that cos(x) is negative on the interval \[(\frac{\pi}{2}+2\pi,\frac{3\pi}{2}+2\pi)\] so cos(2x) would be negative on the interval \[(\frac{1}{2}(\frac{\pi}{2}+2\pi),\frac{1}{2}(\frac{3\pi}{2}+2\pi))\] so in general cos(2x) is negative on the interval(s): \[(\frac{1}{2}(\frac{\pi}{2}+2n \pi)),\frac{1}{2}(\frac{3\pi}{2}+2npi))\] if i didn't make a mistake

myininaya (myininaya):

so \[\cos(2x) \le 0\] when \[x=\frac{npi}{2} or x \in (\frac{1}{2}(\frac{\pi}{2}+2n \pi),\frac{1}{2}(\frac{3\pi}{2}+2n \pi))\]

myininaya (myininaya):

\[x=\frac{n \pi}{2} or x \in (\frac{1}{2}(\frac{\pi}{2}+2n \pi),\frac{1}{2}(\frac{3\pi}{2}+2n \pi))\]

OpenStudy (anonymous):

Im confused so cos = \[\sqrt{2}\div2\] and i have to jus all reference angles?

myininaya (myininaya):

what?

OpenStudy (anonymous):

i get it thanks

OpenStudy (anonymous):

could you help me with find the inverse of F(x)=x^3+9x7

myininaya (myininaya):

is that 7th power?

OpenStudy (anonymous):

\[x ^{3}+9x+7\]